# Sparse Bayesian multidimensional scaling(s)

### v1

@@ A.3 Completing the Proof
$\mathcal{A}_{N}=\left\{\forall n\in[N],k\in[K],\,p(\{u\,:\,\min(|u-x_{n}|,|u+x_{n}|)\leq c_{1}\epsilon_{N}\}|\{\delta_{ny}\}_{y\in G_{n,N}^{(k)}})\geq 1-e^{-c_{2}\ell_{N}\epsilon_{N}^{2}}\right\}$ (26)
@@ A.4 Extending Theorem 1 to Higher Dimensions
It is natural to ask if Theorem 1 holds in higher dimensions. The answer appears to be “yes,” but the only proofs that we are aware of have at least one of the following two substantial flaws: they are noticeably longer or give constants $C$ that scale very poorly with dimension. We give a quick sketch of an argument that most closely mimics our one-dimensional argument, noting the flaws as we do so.
In the current argument, we invoke Theorem 1 of Wong and Shen, (1995) twice: once in Lemma 1 on the “single row” $G_{n,N}^{(k)}$ to show that we have learned $|x_{n}|$ with high accuracy, and again in Lemma 2 on the “pair of rows with large intersection” $J$ to show that we have learned the sign of $x_{n}$ (as long as $|x_{n}|$ is sufficiently large). To extend this to a higher dimension $D$, we would invoke Theorem 1 of Wong and Shen, (1995) $(D+1)$ times. On the first invocation, we would show that the posterior distribution of $x_{n}$ concentrates near a $(D-1)$-dimensional set that contains the true point. For $1\leq d\leq D$, in the $d$’th invocation, we would show that we have learned that $x_{n}$ is on a certain subset of dimension $(D-d)$ with high accuracy by looking at $d$ rows of the matrix. Thus, after $D$ invocations, we would have shown that $x_{n}$ is recoverable up to a set of dimension 0.
The last invocation would be used to deal with ambiguity on a finite set, as in the one-dimensional case. Most of the required changes would be routine (e.g., in part 2 of Assumption 1, we would need intersections of (D+1) parts of the partition to support our (D+1) invocations of Theorem 1 of Wong and Shen, (1995)). The important flaws come from extending Lemma 3. This calculation is what describes a quantitative sort of identifiability for the model. To extend our arguments to higher dimensions, we need a result along the lines of: “the set of latent points that (i) lie in a set of dimension $(D-d+1)$ and (ii) have a given expected distance will lie in a reasonably nice set of dimension $(D-d)$.” One can easily check this in the initial case $d=1$ when we ignore truncations: for $r>0$, the set of points $x$ such that $E[\delta_{nn^{\prime}}|x_{n}=x]=r$ is a sphere. For fixed $d>1$, the various truncations and conditionings involved in repeatedly using this calculation will result in repeated application of unions, intersections and truncation operations to these spheres. We see no easy way to do quick calculations on the resulting set, and no way at all to obtain estimates with a good dependence on $D$.
@@ Appendix B Additional Plots
Figures B.1 and B.2 are analogous to Figures 1 and 5 from Section 3, but under the sparse model using landmarks (L-sBMDS). Figure B.1 demonstrates that very few landmarks are necessary to achieve high accuracy relative to the number of data points. Figure B.2 plots the raw speed-ups, varying the number of landmarks as the number of data points increases. Finally, Figure B.3 illustrates the posterior distribution of the strain-specific diffusion rates under the B-sBMDS/HMC model using 100 and 200 bands. When the number of bands is 200, we see no apparent difference from the full BMDS plot in Figure 7.
Figure B.1: The mean of the mean squared error (MSE) across all distances using $1$ to $10$ landmarks for $10$ data points and $1$ to $20$ landmarks for $100$ and $1{,}000$ data points. We estimate distances from the inferred locations obtained using an adaptive Hamiltonain Monte Carlo algorithm under landmark sparse Bayesian multidimensional scaling (L-sBMDS). $\sigma^{2(true)}$ is the variance component of the truncated normal noise centered at 00 added to the “true” distance matrix such that $\sigma^{(true)}$ corresponds to the BMDS error standard deviation $\sigma$

### v3

@@ A.3 Completing the Proof
$\mathcal{A}_{N}=\left\{\forall n\in[N],k\in[K],\,p(\{u\,:\,\min(|u-x_{n}|,|u+x_{n}|)\leq c_{1}\epsilon_{N}\}|\{\delta_{ny}\}_{y\in G_{n,N}^{(k(n))}})\geq 1-e^{-c_{2}\ell_{N}\epsilon_{N}^{2}}\right\}$ (26)
@@ A.4 Extending Theorem 1 to Higher Dimensions
It is natural to ask if Theorem 1 holds in higher dimensions. The answer appears to be “yes,” but the only proofs that we are aware of have at least one of the following two substantial flaws: they are noticeably longer or give constants $C$ that scale very poorly with dimension. We give here a quick sketch of a proof that closely mimics our one-dimensional argument. It requires no new ideas, but gives bounds that scale very poorly with respect to dimension.
In our proof of Theorem 1, we invoke Theorem 1 of Wong and Shen, (1995) twice: once in Lemma 1 on the “single row” $G_{n,N}^{(k(n))}$ to show that we have learned $|x_{n}|$ with high accuracy, and again in Lemma 2 on the “pair of rows with large intersection” $J$ to show that we have learned the sign of $x_{n}$ (as long as $|x_{n}|$ is sufficiently large). To extend this to a higher dimension $D$, we would invoke Theorem 1 of Wong and Shen, (1995) $(D+1)$ times. On the first invocation, we would show that the posterior distribution of $x_{n}$ concentrates near a $(D-1)$-dimensional set that contains the true point. For $1\leq d\leq D$, in the $d$’th invocation, we would show that we have learned that $x_{n}$ is on a certain subset of dimension $(D-d)$ with high accuracy by looking at $d$ rows of the matrix. Thus, after $D$ invocations, we would have shown that $x_{n}$ is recoverable up to a set of dimension 0. These calculations are nearly identical to the calculations in the current proof.
The last invocation would be used to deal with ambiguity on a finite set, as in the one-dimensional case. Most of the required changes would be routine (e.g., in part 2 of Assumption 1, we would need intersections of (D+1) parts of the partition to support our (D+1) invocations of Theorem 1 of Wong and Shen, (1995)). The biggest change comes in proving the natural analogue to Lemma 3. This calculation is what describes a quantitative sort of identifiability for the model. To extend our arguments to higher dimensions, we need a result along the lines of: “the set of latent points that (i) lie in a set of dimension $(D-d+1)$ and (ii) have a given expected distance $r$ will lie in a reasonably nice set of dimension $(D-d)$.” Ignoring truncations, one can easily check see that the following is true in dimension $d=1$: for $r>0$, the set of points $x$ such that $E[\delta_{nn^{\prime}}|x_{n}=x]=r$ is a sphere. When we allow for truncations in dimension $d=1$, we merely need to slice off part of the sphere, and so Lemma 3 is straightforward. For fixed $d>1$, the various truncations and conditionings involved in repeatedly using this calculation will result in repeated application of unions, intersections and truncation operations to these spheres. In dimension $d=2$, proving the resulting analogue of Lemma 3 in this way is a straightforward but very messy calculus exercise. Unfortunately, we see no easy way to do quick calculations on the resulting set in arbitrary dimension, and no way at all to obtain estimates with reasonable dependence on $D$.
@@ Appendix B Additional Plots
Figures B.1 and B.2 are analogous to Figures 1 and 7 from Section 3, but under the sparse model using landmarks (L-sBMDS). Figure B.1 demonstrates that very few landmarks are necessary to achieve high accuracy relative to the number of data points. Figure B.2 plots the raw speed-ups, varying the number of landmarks as the number of data points increases. Finally, Figure B.3 illustrates the posterior distribution of the strain-specific diffusion rates under the B-sBMDS/HMC model using 100 and 200 bands. When the number of bands is 200, we see no apparent difference from the full BMDS plot in Figure 9.
Figure B.1: The mean of the mean squared error (MSE) across all distances using $1$ to $10$ landmarks for $10$ data points and $1$ to $20$ landmarks for $100$ and $1{,}000$ data points. We estimate Euclidean distances from the inferred locations obtained using an adaptive Hamiltonian Monte Carlo algorithm under landmark sparse Bayesian multidimensional scaling (L-sBMDS). $\sigma^{2}_{true}$ is the variance component of the truncated normal noise centered at 00 added to the “true” distance matrix such that $\sigma_{true}$ corresponds to the BMDS error standard deviation $\sigma$
@@
Figure B.4: Ratio of the mean Hellinger distance for various dimensionalities ($D$) and number of data points ($N)$. Hellinger distances are computed from the posterior distributions of the estimated Euclidean distances between full BMDS and banded sBMDS with $\lceil D\sqrt{N}\rceil$ bands. Data is generated under Gaussian assumptions as described in Section 3. For fixed $D$, we calculate the ratio as the mean Hellinger distance with $N$ data points and with $N=100$. The near horizontal lines across the underlying dimensionality indicate that the selected number of bands obtains results that are very similar to the full model despite the increase in dimensions. The ratio of the Hellinger distance is smaller for larger $N$ as error goes down at some statistical rate, e.g., of $\frac{1}{\sqrt{N}}$.